If a simple pendulum is taken to a height equal to the radius of the earth, how many times will the period of oscillation increase?
1
Expert's answer
2021-09-20T10:00:33-0400

The period of oscillation of a simple pendulum is given by

"T_0=2\pi\sqrt{\frac{l_0}{g_0}}"

The acceleration due gravity

"g_0=G\frac{M}{R^2}"

At the height "H=R" above Earth's surface

"g=G\frac{M}{(2R)^2}=\frac{1}{4}g_0"

Hence

"T=2\pi\sqrt{\frac{l_0}{g}}=2T_0"

Answer: two times.